The function f(x, y) =cos √x^3 +y^3 is a homogenous function. Say true or false.

1
Expert's answer
2022-01-18T06:54:48-0500
"f(x, y)=\cos \sqrt{x^3+y^3}"

"f(kx, ky)=\cos \sqrt{(kx)^3+(ky)^3}"

"=\cos (k^{3/2}\sqrt{x^3+y^3})"

Therefore the statement that the function "f(x, y)=\cos \sqrt{x^3+y^3}" is a homogenous function is False.